(FIRST QUESTION)
* Peter has three sticks measuring 19 cm,23 cm and 27cm.He lays them down to form a triangle.
(1)What is the measure of the angle formed by the 19-cm and 23-cm sides to the nearest degree?
(2)What is the measure of the angle formed by the 19-cm and the 27-cm sides to the nearest degree?
(SECOND QUESTION)
*A ladder 8 meters long leans against the wall of the building.If the foot of the ladder makes an angle of 68 degree with the ground,how far is the base of the ladder from the wall?
(3RD QUESTION)
*A newly-wed couple wants to purchase a piece of land which is triangular in shape.If the two sides of the lot measure 25 m and 30m and their included angle is 60 degree,what is it's area? If the lot costs Php 2,000 per square meter,how much would it cost the couple to buy the lot?




I need help.Please I need solution.also.thankyou~ v:)

Answer by stanbon(75887)

Peter has three sticks measuring 19 cm,23 cm and 27cm.He lays them down to form a triangle.
Use the Law of Cosines:: a^2 = b^2+c^2 - 2bc*cos(A)
----------
(1)What is the measure of the angle formed by the 19-cm and 23-cm sides to the nearest degree?
Let that angle be "A"::
cos(A) = [27^2-19^2-23^2)/(2*19*23 = -0.1842
A = arccos(-0.1842) = 100.61 degrees
-----------------------------------------
(2)What is the measure of the angle formed by the 19-cm and the 27-cm sides to the nearest degree?
I'll leave that to you.
-------------------------
(SECOND QUESTION)
*A ladder 8 meters long leans against the wall of the building.If the foot of the ladder makes an angle of 68 degree with the ground,how far is the base of the ladder from the wall?
Sketch the picture::
You have a right triangle with base angle = 68 degree and hypotenuse = 8 m
cos(68) = base/8 m
---
base = 8(cos(68) = 3 meters
----------------------------------
(3RD QUESTION)
*A newly-wed couple wants to purchase a piece of land which is triangular in shape.If the two sides of the lot measure 25 m and 30m and their included angle is 60 degree,what is it's area? If the lot costs Php 2,000 per square meter,how much would it cost the couple to buy the lot?
Area = a*b*cos(C)
Area = 25*30*cos(60) = 375 sq meters
----
Cost = 2000*375 Php
---------------
Cheers,
Stan H.
-------------